For the following set of cash flows, Year Cash Flow

0 –$7,000

1 6,500

2 3,700

3 3,100

a. What is the NPV at a discount rate of 0 percent?

b. What is the NPV at a discount rate of 11 percent?

c. What is the NPV at a discount rate of 19 percent?

d. What is the NPV at a discount rate of 24 percent?
Answer #1

a.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=6,500+3,700+3,100

=13300

NPV=Present value of inflows-Present value of outflows

=13300-7000

=$6300

b.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=6,500/1.11+3,700/1.11^2+3,100/1.11^3

=11125.55

NPV=Present value of inflows-Present value of outflows

=11125.55-7000

=$4125.55(Approx).

c.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=6,500/1.19+3,700/1.19^2+3,100/1.19^3

=9914.58

NPV=Present value of inflows-Present value of outflows

=9914.58-7000

=$2914.58(Approx).

d.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=6,500/1.24+3,700/1.24^2+3,100/1.24^3

=9274.19

NPV=Present value of inflows-Present value of outflows

=9274.19-7000

=$2274.19(Approx).